OpenStudy (anonymous):

If 44 percent of the seventh grade class are girls, how many of the 150 seventh grade students are girls? A. 3 B. 44 C. 66 D. 84

OpenStudy (campbell_st):

try 44/100 x 150 =

OpenStudy (mathstudent55):

You need to find 44% of 150 In math, "of" usually means times. Also, % means 1/100 or 0.01, so 44% of 150 means 44 * 1/100 * 150

OpenStudy (anonymous):

150*.44=???

OpenStudy (campbell_st):

thats correct

OpenStudy (mayankdevnani):

@nicc96 when you simplify it... \[\huge \frac{44}{100} \times 150=????\]

OpenStudy (anonymous):

you multiply 150 with .44 because the .44 represents the 44% of girls. The 150 is the total seventh grade class. to find out 44% you in turn multiply. 150*.44= your answer.

OpenStudy (mayankdevnani):

\[\huge \frac{44}{100} \times 150=66\] so correct option is C

OpenStudy (anonymous):

thx for help

OpenStudy (mayankdevnani):

welcome

OpenStudy (mayankdevnani):

@nicc96 any confusion????